OpenStudy (anonymous):

Part 1: Explain, in complete sentences, which method you would use to solve the following system of equations. Part 2: Explain why you chose that method Part 3: Provide the solution to the system. x – 2y + z = 0 2x – 3y – 4z = –9 x + 2y – 5z = 0

OpenStudy (anonymous):

I would multiply x –2y + z = 0 by -1, then add x –2y + z = 0 to x + 2y – 5z = 0 (-x + 2y - z = 0)+(x + 2y -5z = 0) 4y -6z = 0 Next, multiply x + 2y - 5z = 0 by -2 and add that to 2x - 3y - 4z = -9 (-2x - 4y + 10z = 0)+(2x - 3y - 4z = -9) -7y + 6z = -9 Let's add these two equations. (-7y + 6z = -9)+(4y -6z = 0) -3y = -9 y = 3 So, now we have a y-value. Let's plug it into our 3 original equations. x – 2y + z = 0 2x – 3y – 4z = –9 x + 2y – 5z = 0 x - 2(3) + z = 0 2x - 3(3) - 4z = -9 x + 2(3) - 5z = 0 x + z = 6 2x - 4z = 0 x -5z = -6 Now, let's solve for x in two of the equations. x = 6-z x = -6 +5z Set these equal to each other. 6-z = -6 + 5z 12 = 6z z = 2 Now that we have a z-value, we can plug in z in these three equations. x + z = 6 2x - 4z = 0 x -5z = -6 x + (2) = 6 2x - 4(2) = 0 x - 5(2) = -6 x = 4 So, our solution to this system of equations is: (4, 3, 2) assuming (x, y, z)

OpenStudy (anonymous):

thanks that is what i got!

OpenStudy (anonymous):

:) Good job.

OpenStudy (anonymous):

check out the one i just posted PLEASE! :(